# Semiclassical Canovaccio for Composite Operators

### v1

@@ 5.5.2 $O(N)$ CFT
Moving to the general $\mathcal{O}(N)$ model, the ground state energy is obtained setting $q_{1,\ell}=q_{2,\ell}=0$ in Eq. (5.5) yielding
$\Delta_{n}=n\left(1-\frac{\epsilon}{2}\right)+\frac{(3n+N-4)n\epsilon}{2(N+8)}+\frac{n^{2}\epsilon^{2}(-17n(N+8)+N(10-11N)+604)}{4(N+8)^{3}}+\mathcal{O}(\epsilon^{2}n,\epsilon^{3})\,$, (122)
which corresponds to the singlet operator $(\phi_{a}\phi_{a})^{n/2}$. In the remaining part of this section, we will simply write $\phi^{n}$ as a shorthand for the $\mathcal{O}(N)$ singlet $(\phi_{a}\phi_{a})^{n/2}$ where $n$ has to be even.
@@
$\Delta_{n}^{\text{3-loop}}=n+P_{1}(n)\epsilon+P_{2}(n)\epsilon^{2}+P_{3}(n)\epsilon^{3}+\mathcal{O}(\epsilon^{4})\ ,$ (123)
$P_{3}(n)=\frac{n}{16(N+8)^{5}}\Bigg[375n^{3}+\frac{(32(N+8)(N+26)\zeta_{3}+3N(95N-312)-21648)n^{2}}{(N+8)}$
@@
We now excite $q_{1,\ell}$ modes while keeping $q_{2,\ell}=0$. Notice from Eq. (113) that $\nu_{1,1}\neq{\cal T}$, so that exciting $\ell=1$ modes does not lead to the descendant states. This reflects the fact that for $N>1$ there are primary operators of the form $\Box\phi^{n}$ and $\partial\phi^{n}$ which are absent for $N=1$. Let us start by discussing the effect of these modes.
@@
    $+\mathcal{O}(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3})\,.$ (125)
    The related tensor decomposition informs us that the families of operators with this scaling dimension are $\partial^{2}\phi^{n}$ and $\Box\phi^{n}$. Indeed, the above matches the known Henriksson:2022rnm $1$-loop scaling dimension of the Lorentz scalar, $\mathcal{O}(N)$ singlet operators $\Box\phi^{n},\ n=4,6,8,$... which we report in Eq. (170).
@@ 6 The $O(N)$ $\phi^{6}$ theory in $d=3-\epsilon$
$\frac{\lambda_{*}^{2}}{4\pi^{2}}=\frac{\epsilon}{3N+22}+\mathcal{O}({\epsilon^{2}})\ .$ (142)
@@
$\begin{split}\Delta_{n}&=\frac{n}{2}(1-\epsilon)+\frac{n(n-2)(5n+3N-8)\epsilon}{6(3N+22)}+\frac{\epsilon^{2}}{192(3N+22)^{3}}\Big(n(8(-131n^{4}(22+3N)\\ &-2n^{3}(22+3N)(-576+71N)+n^{2}(-66664+3N(4796+(802-35N)N))\\ &+n(64272+N(-36020-996N+669N^{2}))-2(7856+N(-10308+N(1792\\ &+435N))))-(-2+n)(118496+27n^{2}(24+N)(22+3N)+N(-26536+N(-7448\\ &+N(-140+3N)))+2n(-47656+N(-3586+N(595+22N))))\pi^{2})\Big)+\mathcal{O}(\epsilon^{3})\ .\end{split}$ (143)
@@ 6.1 Classical solution and leading order: $C_{0}$
$\begin{split}a&=(2\mu)^{\frac{1}{2}}\left(\frac{2\left(\sqrt{(m-1)m+1}-1\right)}{36g^{2}(m-1)^{2}\sqrt{(m-1)m+1}}\right.\\ &+\left.\frac{m\left(-2\sqrt{(m-1)m+1}+m\left(-2m+2\sqrt{(m-1)m+1}+3\right)+3\right)}{36g^{2}(m-1)^{2}\sqrt{(m-1)m+1}}\right)^{1/4}\ ,\end{split}$ (146)
@@
$I$ $=16\pi\omega a^{2}\int\frac{\sqrt{1-u^{2}}\sqrt{m\left(u^{2}-1\right)+1}}{\left(bu^{2}+1\right)^{3}}\,du$
$=-\frac{16\pi\omega a^{2}}{8b^{2}(b+1)\sqrt{1-m}(b(m-1)+m)}\Bigg[\left((3b+4)b^{3}\right$.
@@
$C_{0}=\Omega_{d-1}\braket{T_{00}}\rvert_{d=3}=\frac{\sqrt{2}\pi\omega^{2}}{3gn}\sqrt{1-m+m^{2}-2(m-2)(m+1)(2m-1)\omega^{2}}\ .$ (151)
Using the solution (150) of the Bohr-Sommerfeld constraint and reverting back to $\lambda=ig$ and expanding for $\lambda n\ll 1$, we get
$C_{0}=\frac{1}{2}+\frac{5}{24\pi^{2}}(\lambda n)^{2}-\frac{131}{384\pi^{4}}(\lambda n)^{4}+\frac{4915}{4608\pi^{6}}(\lambda n)^{6}+\mathcal{O}\left({(\lambda n)^{9}}\right)\ ,$ (152)
@@
in agreement with the generic expectation $\Delta_{n,q_{\ell}}\underset{n\to\infty}{\sim}n^{\frac{d}{d-1}}$. Despite appearances, the expression above is real for all values of $\lambda$ and $n$.
@@ 6.2 Leading quantum correction: $C_{1}$
$C_{1}=\frac{1}{2{\cal T}}\sum^{\infty}_{\ell=1}\left(n_{\ell}\Big[\nu_{||,\ell}+(N-1)\nu_{\perp,\ell}\Big]+2q_{||,\ell}\nu_{||,\ell}+2q_{\perp,\ell}\nu_{\perp,\ell}\right)\,$, (158)
@@
$\begin{split}\frac{\nu_{||,\ell}}{{\cal T}}&=\ell-\frac{5(\ell-1)}{(2\ell+1)(3N+22)}n^{2}\epsilon\\ &+\frac{5(\ell(\ell+1)(\ell(2\ell(4\ell(\ell(262\ell+325)-267)-1447)-647)+936)+90)}{12\ell(2\ell+1)^{3}\left(4(\ell+2)\ell^{2}+\ell-3\right)(3N+22)^{2}}n^{4}\epsilon^{2}+\mathcal{O}\left({\epsilon^{3}n^{6}}\right)\ ,\end{split}$ (161)
$\begin{split}\frac{\nu_{\perp,\ell}}{{\cal T}}&=\ell-\frac{(5\ell+1)}{(2\ell+1)(3N+22)}n^{2}\epsilon\\ &+\frac{(5\ell(\ell+1)(\ell(2\ell(4\ell(\ell(262\ell+589)+261)-415)-695)-72)+18)}{12\ell(2\ell+1)^{3}\left(4(\ell+2)\ell^{2}+\ell-3\right)(3N+22)^{2}}n^{4}\epsilon^{2}+\mathcal{O}\left({\epsilon^{3}n^{6}}\right)\ ,\end{split}$ (162)
@@
    $\gamma_{n,\delta_{\ell,2}}=\frac{(n-3)(n-2)(5n+4)}{150}\epsilon+\mathcal{O}(\epsilon^{2})\,,\qquad n\geq 2\,$, (164)
@@
    $\gamma_{n,\delta_{\ell,3}}=\frac{(n-3)\left(7n^{2}-12n-40\right)}{210}\epsilon+\mathcal{O}(\epsilon^{2})\,,\qquad n\geq 3\,$, (165)
@@
    $\gamma_{n,2\delta_{\ell,2}}=\frac{(n-4)(n-2)(5n+3)}{150}\epsilon+\mathcal{O}(\epsilon^{2})\,,\qquad n\geq 4\,$, (166)
@@
    $\gamma_{(\pm)}=\frac{35n^{3}-182n^{2}+32n+684\pm\sqrt{49n^{4}+812n^{3}+37228n^{2}-287184n+508176}}{1050}\epsilon+\mathcal{O}(\epsilon^{2})\,$, (167)
@@
    $\gamma_{n,\delta_{\ell,2}+\delta_{\ell,3}}=\frac{35n^{3}-207n^{2}+64n+600}{1050}\epsilon+\mathcal{O}(\epsilon^{2})\,,\qquad n\geq 5\,$, (168)
@@ Appendix A Review of perturbative results in the $\phi^{4}$ model
We collect here the known perturbative results for the anomalous dimension of composite operators in the $\phi^{4}$ theory that we use to compare with the semiclassical calculations presented in Sec. 5.5. These results have been determined in Henriksson:2022rnm ; Henriksson:2025vyi ; Kehrein:1994ff ; Kehrein:1992fn ; Derkachov:1997gc . For the reader’s convenience, we also recall the values of $q_{1,\ell}$ and $q_{2,\ell}$ used to construct the operators from the semiclassical methodology. First, for Lorentz scalar $\mathcal{O}(N)$ singlet operators of the form $\phi^{n},\ n\geq 1$, we have
@@
$q_{1,\ell}=2\delta_{\ell,1},\ q_{2,\ell}=0:\qquad\gamma_{n,\{2\delta_{\ell,1}\}}=\frac{3n^{2}+(N-12)n+8}{2(N+8)}\epsilon+\mathcal{O}(\epsilon^{2})\ ,$ (170)
@@
0 $\phi^{n},\ n\geq 1$ $\frac{n(n-1)}{6}\epsilon-\frac{n(17n^{2}-67n+47)}{324}\epsilon^{2}+\mathcal{O}(\epsilon^{3})$
$\delta_{\ell,2}$ $\partial^{2}\phi^{n},\ n\geq 2$ $\frac{(n-2)(3n+1)}{18}\epsilon+\mathcal{O}(\epsilon^{2})$
$\delta_{\ell,3}$ $\partial^{3}\phi^{n},\ n\geq 3$ $\frac{n^{2}-2n-2}{6}\epsilon+\mathcal{O}(\epsilon^{2})$
$\delta_{\ell,4}$ $\partial^{4}\phi^{n},\ n\geq 2$ $\frac{(n-2)(5n-1)}{30}\epsilon+\mathcal{O}(\epsilon^{2})$
$\delta_{\ell,5}$ $\partial^{5}\phi^{n},\ n\geq 3$ $\frac{3n^{2}-7n-2}{18}\epsilon+\mathcal{O}(\epsilon^{2})$
$2\delta_{\ell,2}$ $\partial^{4}\phi^{n},\ n\geq 4$ $\frac{3n^{2}-7n-12}{18}\epsilon+\mathcal{O}(\epsilon^{2})$
$\partial^{2}\Box\phi^{n},\ n\geq 4$ $\frac{3n^{2}-7n-4}{18}\epsilon+\mathcal{O}(\epsilon^{2})$
$\Box^{2}\phi^{n},\ n\geq 4$ $\frac{n(3n-7)}{18}\epsilon+\mathcal{O}(\epsilon^{2})$
$\delta_{\ell,2}+\delta_{\ell,3}$ $\partial^{5}\phi^{n},\ n\geq 5$ $\frac{3n^{2}-8n-20}{18}\epsilon+\mathcal{O}(\epsilon^{2})$
$\partial^{3}\Box\phi^{n},\ n\geq 5$ $\frac{3n^{2}-8n-10}{18}\epsilon+\mathcal{O}(\epsilon^{2})$
$\partial\Box^{2}\phi^{n},\ n\geq 5$ $\frac{3n^{2}-8n-4}{18}\epsilon+\mathcal{O}(\epsilon^{2})$

### v2

@@ 5.5.2 $O(N)$ CFT
Moving to the general $O(N)$ model, the ground state energy is obtained setting $q_{1,\ell}=q_{2,\ell}=0$ in Eq. (5.5) yielding
$\Delta_{n}=n\left(1-\frac{\epsilon}{2}\right)+\frac{(3n+N-4)n\epsilon}{2(N+8)}+\frac{n^{2}\epsilon^{2}(-17n(N+8)+N(10-11N)+604)}{4(N+8)^{3}}+\mathcal{O}\left(\epsilon^{2}n,\epsilon^{3}\right)\,$, (123)
which corresponds to the singlet operator $(\phi_{a}\phi_{a})^{n/2}$. In the remaining part of this section, we will simply write $\phi^{n}$ as a shorthand for the $O(N)$ singlet $(\phi_{a}\phi_{a})^{n/2}$ where $n$ has to be even.
@@
$\Delta_{n}^{\text{3-loop}}=n+P_{1}(n)\epsilon+P_{2}(n)\epsilon^{2}+P_{3}(n)\epsilon^{3}+\mathcal{O}\left(\epsilon^{4}\right)\ ,$ (124)
$P_{3}(n)=\frac{n}{16(N+8)^{3}}\Bigg[375n^{3}+\frac{(32(N+8)(N+26)\zeta_{3}+3N(95N-312)-21648)n^{2}}{(N+8)}$
@@
We now excite $q_{1,\ell}$ modes while keeping $q_{2,\ell}=0$. Notice from Eq. (5.5) that $\nu_{1,1}\neq{\cal T}$, so that exciting $\ell=1$ modes does not lead to the descendant states. This reflects the fact that for $N>1$ there are primary operators of the form $\Box\phi^{n}$ and $\partial\phi^{n}$ which are absent for $N=1$. Let us start by discussing the effect of these modes.
@@
    $+\mathcal{O}\left(\epsilon n^{0},\epsilon^{2}n,\epsilon^{3}\right)\,.$ (126)
    The related tensor decomposition informs us that the families of operators with this scaling dimension are $\partial^{2}\phi^{n}$ and $\Box\phi^{n}$. Indeed, the above matches the known Henriksson:2022rnm $1$-loop scaling dimension of the Lorentz scalar, $O(N)$ singlet operators $\Box\phi^{n},\ n=4,6,8,$... which we report in Eq. (171).
@@ 6 The $O(N)$ $\phi^{6}$ theory in $d=3-\epsilon$
$\frac{\lambda_{*}^{2}}{4\pi^{2}}=\frac{\epsilon}{3N+22}+\mathcal{O}\left(\epsilon^{2}\right)\ .$ (143)
@@
$\Delta_{n}$ $=\frac{n}{2}(1-\epsilon)+\frac{n(n-2)(5n+3N-8)\epsilon}{6(3N+22)}+\frac{\epsilon^{2}}{192(3N+22)^{3}}\bigg(-1048n^{5}\Big(3N+22\Big)$
$+n^{4}\Big(-16\Big(71N-576\Big)\Big(3N+22\Big)-27\pi^{2}\Big(N+24\Big)\Big(3N+22\Big)\Big)$
$+n^{3}\Big(54\pi^{2}\Big(N+24\Big)\Big(3N+22\Big)+8\Big(3N\Big(\Big(802-35N\Big)N+4796\Big)-66664\Big)$
$-2\pi^{2}\Big(N\Big(N\Big(22N+595\Big)-3586\Big)-47656\Big)\Big)+n^{2}\Big(8\Big(N\Big(669N^{2}-996N-36020\Big)$
$+64272\Big)-\pi^{2}N\Big(N\Big(N\Big(3N-140\Big)-7448\Big)-26536\Big)+4\pi^{2}\Big(N\Big(N\Big(22N+595\Big)$
$-3586\Big)-47656\Big)-118496\pi^{2}\Big)+n\Big(2\pi^{2}N\Big(N\Big(N\Big(3N-140\Big)-7448\Big)-26536\Big)$
$-16\Big(N\Big(N\Big(435N+1792\Big)-10308\Big)+7856\Big)+236992\pi^{2}\Big)\bigg)+\mathcal{O}(\epsilon^{3})$ (144)
@@ 6.1 Classical solution and leading order: $C_{0}$
$\begin{split}a&=(2\mu)^{\frac{1}{2}}\left(\frac{2(m-1)m-\frac{(m-2)(m+1)(2m-1)}{\sqrt{(m-1)m+1}}+2}{36g^{2}(m-1)^{2}}\right)^{1/4}\ ,\end{split}$ (147)
@@
$I$ $=16\pi r^{2}\omega a^{2}\int\frac{\sqrt{1-u^{2}}\sqrt{m\left(u^{2}-1\right)+1}}{\left(bu^{2}+1\right)^{3}}\,du$
$=-\frac{16\pi r^{2}\omega a^{2}}{8b^{2}(b+1)\sqrt{1-m}(b(m-1)+m)}\Bigg[\left((3b+4)b^{3}\right$.
@@
$C_{0}=\Omega_{d-1}r^{d}\braket{T_{00}}\rvert_{d=3}=\frac{\sqrt{2}\pi(r\omega)^{2}}{3gn}\sqrt{1-m+m^{2}-2(m-2)(m+1)(2m-1)(r\omega)^{2}}\ .$ (152)
Using the solution (151) of the Bohr-Sommerfeld constraint, reverting back to $\lambda=ig$, and expanding for $\lambda n\ll 1$, we get
$C_{0}=\frac{1}{2}+\frac{5}{24\pi^{2}}(\lambda n)^{2}-\frac{131}{384\pi^{4}}(\lambda n)^{4}+\frac{4915}{4608\pi^{6}}(\lambda n)^{6}+\mathcal{O}\left({(\lambda n)^{8}}\right)\ ,$ (153)
@@
in agreement with the generic expectation $\Delta_{n,q_{\ell}}\underset{n\to\infty}{\sim}n^{\frac{d}{d-1}}$. Despite appearances, the expression above is real for all the real values of $\lambda$ and $n$.
@@ 6.2 Leading quantum correction: $C_{1}$
$C_{1}=\frac{r}{2{\cal T}}\sum^{\infty}_{\ell=1}\left(n_{\ell}\Big[\nu_{||,\ell}+(N-1)\nu_{\perp,\ell}\Big]+2q_{||,\ell}\nu_{||,\ell}+2q_{\perp,\ell}\nu_{\perp,\ell}\right)\,$, (159)
@@
$\begin{split}\frac{r\nu_{||,\ell}}{{\cal T}}&=\ell-\frac{5(\ell-1)}{(2\ell+1)(3N+22)}n^{2}\epsilon\\ &+\frac{5(\ell(\ell+1)(\ell(2\ell(4\ell(\ell(262\ell+325)-267)-1447)-647)+936)+90)}{12\ell(2\ell+1)^{3}\left(4(\ell+2)\ell^{2}+\ell-3\right)(3N+22)^{2}}n^{4}\epsilon^{2}+\mathcal{O}\left({\epsilon^{3}n^{6}}\right)\ ,\end{split}$ (162)
$\begin{split}\frac{r\nu_{\perp,\ell}}{{\cal T}}&=\ell-\frac{(5\ell+1)}{(2\ell+1)(3N+22)}n^{2}\epsilon\\ &+\frac{(5\ell(\ell+1)(\ell(2\ell(4\ell(\ell(262\ell+589)+261)-415)-695)-72)+18)}{12\ell(2\ell+1)^{3}\left(4(\ell+2)\ell^{2}+\ell-3\right)(3N+22)^{2}}n^{4}\epsilon^{2}+\mathcal{O}\left({\epsilon^{3}n^{6}}\right)\ ,\end{split}$ (163)
@@
    $\gamma_{n,\delta_{\ell,2}}=\frac{(n-3)(n-2)(5n+4)}{150}\epsilon+\mathcal{O}\left(\epsilon^{2}\right)\,,\qquad n\geq 2\,$, (165)
@@
    $\gamma_{n,\delta_{\ell,3}}=\frac{(n-3)\left(7n^{2}-12n-40\right)}{210}\epsilon+\mathcal{O}\left(\epsilon^{2}\right)\,,\qquad n\geq 3\,$, (166)
@@
    $\gamma_{n,2\delta_{\ell,2}}=\frac{(n-4)(n-2)(5n+3)}{150}\epsilon+\mathcal{O}\left(\epsilon^{2}\right)\,,\qquad n\geq 4\,$, (167)
@@
    $\gamma_{(\pm)}=\frac{35n^{3}-182n^{2}+32n+684\pm\sqrt{49n^{4}+812n^{3}+37228n^{2}-287184n+508176}}{1050}\epsilon+\mathcal{O}\left(\epsilon^{2}\right)\,$, (168)
@@
    $\gamma_{n,\delta_{\ell,2}+\delta_{\ell,3}}=\frac{35n^{3}-207n^{2}+64n+600}{1050}\epsilon+\mathcal{O}\left(\epsilon^{2}\right)\,,\qquad n\geq 5\,$, (169)
@@ Appendix A Review of perturbative results in the $\phi^{4}$ model
We collect here the known perturbative results for the anomalous dimension of composite operators in the $\phi^{4}$ theory that we use to compare with the semiclassical calculations presented in Sec. 5.5. These results have been determined in Henriksson:2022rnm ; Henriksson:2025vyi ; Kehrein:1994ff ; Kehrein:1992fn ; Derkachov:1997gc . For the reader’s convenience, we also recall the values of $q_{1,\ell}$ and $q_{2,\ell}$ used to construct the operators from the semiclassical methodology. First, for Lorentz scalar $O(N)$-singlet operators of the form $\phi^{n},\ n\geq 1$, we have
@@
$q_{1,\ell}=2\delta_{\ell,1},\ q_{2,\ell}=0:\qquad\gamma_{n,\{2\delta_{\ell,1}\}}=\frac{3n^{2}+(N-12)n+8}{2(N+8)}\epsilon+\mathcal{O}\left(\epsilon^{2}\right)\ ,$ (171)
@@
0 $\phi^{n},\ n\geq 1$ $\frac{n(n-1)}{6}\epsilon-\frac{n(17n^{2}-67n+47)}{324}\epsilon^{2}+\mathcal{O}\left(\epsilon^{3}\right)$
$\delta_{\ell,2}$ $\partial^{2}\phi^{n},\ n\geq 2$ $\frac{(n-2)(3n+1)}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\delta_{\ell,3}$ $\partial^{3}\phi^{n},\ n\geq 3$ $\frac{n^{2}-2n-2}{6}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\delta_{\ell,4}$ $\partial^{4}\phi^{n},\ n\geq 2$ $\frac{(n-2)(5n-1)}{30}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\delta_{\ell,5}$ $\partial^{5}\phi^{n},\ n\geq 3$ $\frac{3n^{2}-7n-2}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$2\delta_{\ell,2}$ $\partial^{4}\phi^{n},\ n\geq 4$ $\frac{3n^{2}-7n-12}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\partial^{2}\Box\phi^{n},\ n\geq 4$ $\frac{3n^{2}-7n-4}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\Box^{2}\phi^{n},\ n\geq 4$ $\frac{n(3n-7)}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\delta_{\ell,2}+\delta_{\ell,3}$ $\partial^{5}\phi^{n},\ n\geq 5$ $\frac{3n^{2}-8n-20}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\partial^{3}\Box\phi^{n},\ n\geq 5$ $\frac{3n^{2}-8n-10}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$
$\partial\Box^{2}\phi^{n},\ n\geq 5$ $\frac{3n^{2}-8n-4}{18}\epsilon+\mathcal{O}\left({\epsilon^{2}}\right)$